OpenStudy (anonymous):

When Julian got on the bus, 8 people got on and 2 got off. At the next stop, 5 people got on and 3 got off. At Julian's stop, he noticed that 7 people got on, 9 got off, and there were 16 people left. How many passengers were on the bus before Julian got on?

OpenStudy (anonymous):

please help its hard

OpenStudy (johnweldon1993):

Well right now I'm trying to figure out if "At Julian's stop, he noticed that 7 people got on, 9 got off" INCLUDES Julian or if it is actually 10 people got off (counting Julian) weird wording...

OpenStudy (anonymous):

I think its not including himself

OpenStudy (anonymous):

just the people in the bus

OpenStudy (anonymous):

help!

OpenStudy (johnweldon1993):

Hmm..I'm going to have to go with 10 lets think..if there were 10 people on the bus to begin with before he got on And 6 people got on (8 get on and 2 leave = 6) that would be 16 people Then 2 people get on (5 people on...3 leave = 2) that would be 18 people Then at Julians stop....2 people leave (9 people leave and 7 get on = -2) that would bring us back to 16 people left

OpenStudy (johnweldon1993):

Make sense?

OpenStudy (anonymous):

not really?

OpenStudy (johnweldon1993):

Sorry for getting off...but okay....Here I'll show you it in equation form Let 'x' = the number of people before Julian got on.... After the first stop we have + 8 people...and - 2 people so x + 8 - 2 After the second stop we have + 5 people and - 3 people so x + 8 - 2 + 5 - 3 After Julains stop we have + 7 people and - 9 people so x + 8 - 2 + 5 - 3 + 7 - 9 And we know there were 16 people leftover...so x + 8 - 2 + 5 - 3 + 7 - 9 = 16 So now lets solve x + 6 + 5 - 3 + 7 - 9 = 16 x + 11 - 3 + 7 - 9 = 16 x + 8 + 7 - 9 = 16 x + 15 - 9 = 16 x + 6 = 16 x = 10 so there were 10 people on the bus before Julian got on